# Grammar-based Game Description Generation using Large Language Models

### v1

@@ IV-A Grammar-based Game Description Generation
We first introduce the minimal grammar $G[y]\subseteq G$ necessary to compose $y$. $G[y]$ is a subset of the full grammar $G$, where $y\in L(G[y])$ and $\forall{r}\in G[y],y\notin L(G[y]\setminus\{r\})$. We can obtain $G[y]$ by parsing $y$ with $G$ and collecting the rules to derive $y$. Each demonstration consists of $(x^{(i)},\,G[y^{(i)}],\,y^{(i)})$.
@@
In the second stage, the generated $\hat{G}[y]$ is added to the prompt to generate the game description $\hat{y}$, and the $y’$ ’ that maximizes the following probability is selected as $\hat{y}$:
@@ IV-B Grammar-based Iterative Decoding
Our rule decoding stage iteratively applies rule decoding while aiming to ensure that the generated $\hat{G}[y]$ is a subset of the original grammar $G$. In our rule decoding stage, the generation process is multi-step, where each step iteratively improves the grammar $\hat{G}[y]$. The goal of each step is to define the non-terminal symbols that were not defined in the grammar $\hat{G}[y]$ generated in the previous step. We illustrate our rule decoding process in Fig. 6. Initially, $\hat{G}[y]$ is generated using the LLM in the same way as in Eq. (1). From $\hat{G}[y]$, only the valid grammar rules $\hat{G}[y]_{\mathrm{valid}}$ are retained, which are included in the original grammar $G$. Among $\hat{G}[y]_{\mathrm{valid}}$, the set of undefined non-terminal symbols $N_{U}$ is extracted. The undefined non-terminal symbols are those that are used on the right-hand side of rules in $\hat{G}[y]_{\mathrm{valid}}$ but are not defined (i.e., they do not appear on the left-hand side of any rules), and can be automatically extracted easily. The set of grammar rules for the undefined non-terminal symbols $G_{N_{U}}$ is extracted from the original grammar $G$. $G_{N_{U}}$ is added to the prompt, and the LLM selects only the necessary rules from $G_{N_{U}}$ based on the query $x$. Since $G_{N_{U}}$ is extracted from the original grammar $G$, it contains more rules than the minimum required for $y$. The role of the LLM is to extract only the rules $\hat{G}_{N_{U}}=\hat{G}[y]\setminus\hat{G}[y]_{\mathrm{valid}}$ necessary to compose $y$ from $G_{N_{U}}$. The rules that maximize the following probability are selected:
$P_{\mathrm{LLM}}(\hat{G}_{N_{U}}|(x^{(i)},G[y^{(i)}],y^{(i)})_{i=1}^{N},x,\hat{G}[y]_{\mathrm{valid}},G_{N_{U}}).$ (3)
$\hat{G}_{N_{U}}$ is combined with $\hat{G}[y]_{\mathrm{valid}}$ to update $\hat{G}[y]$. This process is repeated until there are no more non-terminal symbols or a predetermined limit of updates is reached.
@@
The game description decoding stage consists of iterative application of game description decoding to generate game descriptions that more accurately adhere to the grammar $\hat{G}[y]$. In each step, it aims to complete the parts of $\hat{y}$ generated in the previous step that cannot be parsed. Initially, $\hat{y}$ is generated using the LLM in the same way as in Eq. (2). An Earley parser [45] is then employed using $\hat{G}[y]$. The Earley parser explores the program from left to right, extracting the longest valid subsequence and the subsequent candidate terminal symbols. Using the generated $\hat{y}$, the Earley parser obtains the valid subsequence $\hat{y}_{\mathrm{valid}}$ and the set of candidate terminal symbols $\Omega_{\mathrm{candidate}}$. The LLM selects the most suitable candidate $\omega$ from $\Omega_{\mathrm{candidate}}$. The $\omega^{\prime}$ that maximizes the following probability is then chosen:
$P_{\mathrm{LLM}}(\omega^{\prime}|(x^{(i)},G[y^{(i)}],y^{(i)})_{i=1}^{N},x,\hat{G}[y],\hat{y}_{\mathrm{valid}},\Omega_{\mathrm{candidate}}).$ (4)
@@ V-A Datasets
The game descriptions generated by Ludii [7] that we use for evaluation are obtained from the publicly available Ludii software [46]. In Ludii, each game is assigned a category (e.g., “Tic-Tac-Toe” is categorized under “Board/Space/Line”, and “Tower of Hanoi” under “Puzzle/Planning”). From the same category, one test example and three demonstration examples are extracted as a single instance. Note that only examples where the token length of the game description $y$ is 300 or less are used. The token length is calculated using the tokenizer of Llama3-8B-Instruct [12]. The dataset is constructed from 100 randomly selected instances from all categories, and the same set of instances is used in all evaluations. Each example consists of a natural language instruction $x$, a Ludii game description $y$, and the minimal grammar $G[y]$ required to construct $y$. The instruction $x$ is composed of the metadata “Description” and “Rules” provided by the Ludii game system (see Fig. 2 for an example). The game description $y$ consists of the extended game description derived from the raw text .lud game description files. The extended version is instantiated with primitive options and rulesets to avoid using the Ludii game system’s meta-language features (definitions, options, rulesets, ranges, constants, etc.). We use this extended version to reduce game-specific representations as much as possible. For more details of the Ludii language, please refer to the Ludii language reference [47]. The grammar $G[y]$ is automatically extracted by the parser using the Lark library [48], which is explained in the next subsection.
V-B Implementation Details
@@
We use a parser built with a Python library called Lark [48] to extract the grammar rules of the Ludii game description. The Ludii grammar in the Lark parser employs the rules listed in the Ludii language reference. In our grammar-based iterative decoding, we set an upper limit on the number of iterations to suppress the number of LLM calls. The iteration limit is set to 20 for rule decoding and 30 for game description decoding. Our prompts for grammar-based iterative decoding are shown in Fig 7.
@@
We use the Llama3-8B-Instruct [12] model as the LLM for our experiments. Llama3 is the latest series of open-source large language models provided by Meta. The Llama3 series includes pre-trained models with 8B and 70B parameters. Compared to other models of similar size, the Llama3 series demonstrates superior performance across various benchmarks. To ensure that our framework is practical with limited computational resources such as local environments, we select the smaller 8B model. Additionally, we use the instruction fine-tuning [49, 50] version to make the model accurately respond to the various instructions used in our framework. Since the Llama3 series models are pre-trained with sequences of up to 8,192 tokens, it is necessary to keep the prompts within this length (context length). To address this, we set the number of demonstration examples to three. We use two NVIDIA RTX 6000 Ada GPUs for our experiments.
V-C Evaluation Metrics
@@
  • Grammar Parsability: The proportion of generated game descriptions that can be parsed by the Lark parser, which is based on the Ludii grammar $G$.
  • System Parsability: The proportion of generated game descriptions that can be read by the Ludii game system. Descriptions that are grammatically correct as per the Lark parser but lack any of the necessary functions for a game can not be read by the Ludii game system.
  • Playability: The proportion of generated game descriptions that require a reasonable number of player actions per trial. A trial is considered successful if it fulfills the specified termination conditions in at least 5 out of 10 attempts. In particular, we consider the game to be playable if, by continuously choosing legal actions, the number of actions necessary to reach a termination condition should range between 2 and 1,800. This metric evaluates whether the generated game has a reasonable length of play experience. For a game description to satisfy the Playability criteria, it also has to satisfy the Grammar Parsability and System Parsability criteria. We use the implementation [46] in the Ludii game system for this metric.
@@
  • ROUGE [51]: This is a metric that measures the degree of match between the generated game description and the ground truth game description. It ranges from 0 to 100, with higher values indicating a greater degree of match. The calculation of this metric does not consider syntactic correctness but instead focuses on the similarity between texts. We use the F1 score of ROUGE-L. This value is calculated for each piece of test data, and the average of all data is reported.
V-D Quantitative Results

### v2

@@ IV-A Grammar-based Game Description Generation
$G[y]$ is the minimal grammar that extracts only the rules necessary to generate $y$ from all the rules of $G$. An example of $G[y]$ for tic-tac-toe is shown in Figure 3. $G[y]$ is a subset of the full grammar $G$, where $y\in L(G[y])$ and $\forall{r}\in G[y],y\notin L(G[y]\setminus\{r\})$. For any rule $r$ in $G[y]$, removing $r$ makes it impossible to generate $y$ ($y\notin L(G[y]\setminus{r})$). $G[y]$ is minimal in the sense that it contains exactly the rules required to generate $y$, with no superfluous rules. A parser based on this minimal grammar can determine grammatically valid subsequences and the set of candidate symbols that follow them from the LLM’s responses. The LLM then generates the rest of the game description based on the subsequences and candidate symbol groups. Intuitively, it is expected that repeating this parsing and generation process will result in a more grammatically accurate game description than the initial response from the LLM. This iterative improvement method is explained in Sec. IV-B.
From the perspective of ICL, providing more context to LLMs is expected to have a positive impact. It is known that providing GDL grammar to LLMs can improve the quality of generated game descriptions, especially in terms of grammatical accuracy [14]. However, the token length of Ludii’s grammar using the Llama3-8B-Instruct tokenizer is 15,442, which is longer than the context length of most LLMs. The straight-forward approach of including all the GDL grammar in the prompt as proposed in LLMGG is not feasible without significant more computation power. The average token length of $G[y]$ using the Llama3-8B-Instruct tokenizer is 1,031, which can be added to the prompt within the context length limit for several games. In this case, each demonstration consists of $(x^{(i)},G[y^{(i)}],y^{(i)})$. $G[y^{(i)}]$ is obtained by parsing $y^{(i)}$ with $G$ and collecting the rules necessary to derive $y^{(i)}$.
@@
In the second stage, the generated $\hat{G}[y]$ is added to the prompt to generate the game description $\hat{y}$, and the $y^{\prime}$ that maximizes the following probability is selected as $\hat{y}$:
@@ IV-B Grammar-based Iterative Decoding
Our rule decoding stage iteratively applies rule decoding while aiming to ensure that the generated $\hat{G}[y]$ is a subset of the original grammar $G$. In our rule decoding stage, the generation process is multi-step, where each step iteratively improves the grammar $\hat{G}[y]$. The goal of each step is to define the non-terminal symbols that were not defined in the grammar $\hat{G}[y]$ generated in the previous step. We illustrate our rule decoding process in Fig. 6. Initially, $\hat{G}[y]$ is generated using the LLM in the same way as in Eq. (1). From $\hat{G}[y]$, only the valid grammar rules $\hat{G}[y]_{\mathrm{valid}}$ are retained, which are included in the original grammar $G$. Among $\hat{G}[y]_{\mathrm{valid}}$, the set of undefined non-terminal symbols $N_{U}$ is extracted. Undefined non-terminal symbols are rules that are used on the right-hand side of rules in $\hat{G}[y]_{\mathrm{valid}}$ but are not defined. They can be automatically and easily extracted by identifying symbols not appearing on the left-hand side of any rules. The set of grammar rules for the undefined non-terminal symbols $G_{N_{U}}$ is extracted from the original grammar $G$. $G_{N_{U}}$ is added to the prompt, and the LLM selects only the necessary rules from $G_{N_{U}}$ based on the query $x$. The right-hand side of the rules in $G_{N_{U}}$ includes options that are unnecessary for constructing $y$. The role of the LLM is to select the minimum necessary choices to construct $y$ from these options, and to predict the minimal grammar $G[y]_{N_{U}}=G[y]\setminus\hat{G}[y]_{\mathrm{valid}}$. The rules that maximize the following probability are selected:
$P_{\mathrm{LLM}}(G^{\prime}[y]_{N_{U}}|(x^{(i)},G[y^{(i)}],y^{(i)})_{i=1}^{N},x,\hat{G}[y]_{\mathrm{valid}},G_{N_{U}}).$ (3)
The generated $\hat{G}[y]_{N_{U}}$ is combined with $\hat{G}[y]_{\mathrm{valid}}$ to update $\hat{G}[y]$. This process is repeated until there are no more non-terminal symbols or a predetermined limit of updates is reached.
@@
The game description decoding stage consists of iterative application of game description decoding to generate game descriptions that more accurately adhere to the grammar $\hat{G}[y]$. In each step, it aims to complete the parts of $\hat{y}$ generated in the previous step that cannot be parsed. Initially, $\hat{y}$ is generated using the LLM in the same way as in Eq. (2). An Earley parser [64] is then employed using $\hat{G}[y]$. The Earley parser explores the program from left to right, extracting the longest valid subsequence and the subsequent candidate terminal symbols. Using the generated $\hat{y}$, the Earley parser obtains the valid subsequence $\hat{y}_{\mathrm{valid}}$ and the set of candidate terminal symbols $\Omega_{\mathrm{candidate}}$. Because multiple options for rules are often generated during the rule decoding stage, the LLM selects the optimal terminal symbol $\omega$ from $\Omega_{\mathrm{candidate}}$. The $\omega$ that maximizes the following probability is then chosen:
$P_{\mathrm{LLM}}(\omega|(x^{(i)},G[y^{(i)}],y^{(i)})_{i=1}^{N},x,\hat{G}[y],\hat{y}_{\mathrm{valid}},\Omega_{\mathrm{candidate}}).$ (4)
@@ V-A Datasets
The game descriptions we use for evaluation are obtained from the publicly available Ludii website [65]. In Ludii, each game is assigned a category (e.g., “Tic-Tac-Toe” is categorized under “Board/Space/Line”, and “Tower of Hanoi” under “Puzzle/Planning”). From the same category, one test example and three demonstration examples are extracted as a single instance, leveraging the known benefit of using semantically similar examples to improve ICL performance [66]. Note that only examples where the token length of the game description $y$ is 300 or less are used. The token length is calculated using the tokenizer of Llama3-8B-Instruct [9]. The evaluation dataset is constructed from 100 randomly selected instances from all categories, and the same set of instances is used in all evaluations. Each example consists of a natural language instruction $x$, a Ludii game description $y$, and the minimal grammar $G[y]$ required to construct $y$. The instruction $x$ is composed of the metadata “Description” and “Rules” provided by the Ludii game system (see Fig. 1 for an example). To improve the generality of the dataset, we use game descriptions where the unique functions defined within each game are expanded as $y$. The extended version is instantiated with primitive options and rulesets to avoid using the Ludii game system’s meta-language features (definitions, options, rulesets, ranges, constants, etc.). A similar process is adopted in [46]. For more details of the Ludii language, please refer to the Ludii language reference [67]. The grammar $G[y]$ is automatically extracted by the parser using the Lark library [68], which is explained in the next subsection.
V-B Methods
We compare the following methods:
  • Random: A baseline method that generates the grammar $\hat{G}[y]$ necessary for constructing the game description $y$. Based on $\hat{G}[y]$, the method randomly samples the next expressions following “(game” to generate the game description.
  • Game Description Generation (GDG): A baseline method that directly predicts the game description $y$ from the demonstration examples and the query $x$ without using the predicted grammar $\hat{G}[y]$.
  • Grammar-based Game Description Generation (GGDG, ours): Our proposed method, which first generates the grammar $\hat{G}[y]$ required to construct the game description $y$ and subsequently generates $\hat{y}$ based on $\hat{G}[y]$. This method employs both rule decoding and game description decoding.
  • SFT+GDG: A baseline method that predicts the game description from a query using an LLM with supervised fine-tuning (SFT). For SFT training data, pairs of query $x$ and game description $y$ are created from games available on the Ludii portal [65] that are not included in the evaluation dataset.
  • SFT+GGDG (ours): A method that combines our proposed GGDG with SFT. For rule decoding, an LLM with SFT applied to query $x$ and grammar $G[y]$ pairs is used. For game description decoding, the same LLM model as in SFT+GDG is used, trained with SFT on query $x$ and game description $y$ pairs.
Since the model acquires knowledge of Ludii through SFT, demonstration examples are not used, i.e., zero-shot inference.
V-C Implementation Details
@@
We use a parser built with a Python library called Lark [68] to extract the grammar rules of the Ludii game description. The Ludii grammar in the Lark parser employs the rules listed in the Ludii language reference. In our grammar-based iterative decoding, we set an upper limit on the number of iterations to suppress the number of LLM calls. Based on the ablation study in Sec. VI-C, the iteration limit is set to 20 for rule decoding and 10 for game description decoding. Our prompts for grammar-based iterative decoding are shown in Fig 7.
@@
We use the Llama3-8B-Instruct [9] model as our LLM. We chose an open-source LLM to ensure research reproducibility. Unlike commercial APIs, our approach is not affected by API changes or model updates. Llama3 is the latest series of open-source LLMs provided by Meta. The Llama3 series includes pre-trained models with 8B and 70B parameters. Compared to other models of similar size, the Llama3 series demonstrates superior performance across various benchmarks. To ensure that our framework is practical in environments with limited computational resources, such as local deployments, we select the smaller 8B model. Since the Llama3 models are pre-trained to accept sequences of up to 8,192 tokens, it is necessary to keep the prompts within this context length. To address this, we set the number of demonstration examples to three. We use two NVIDIA RTX 6000 Ada GPUs for our experiments.
For SFT models, we apply SFT to Llama3-8B-Instruct using low-rank adaptation (LoRA) [69]. We set the SFT sequence length to 4096, LoRA alpha and $r$ to 16, and train the model for 3 epochs with a learning rate of 1e-4 and a warmup ratio of 0.03.
TABLE II: Comparison with baseline methods. The best results are in bold.
[…]
Random 1.3$\pm$0.7 1.0$\pm$0.6 9.8$\pm$0.8 0.97$\pm$0.02
GDG 27.0$\pm$1.2 26.3$\pm$0.7 63.5$\pm$0.6 0.75$\pm$0.01
GGDG (ours) 64.0$\pm$1.5 56.7$\pm$2.3 60.5$\pm$0.6 0.46$\pm$0.03
SFT+GDG 59.7$\pm$2.4 58.0$\pm$0.6 64.0$\pm$0.5 0.44$\pm$0.01
SFT+GGDG (ours) 72.0$\pm$2.1 70.3$\pm$1.8 63.8$\pm$0.7 0.33$\pm$0.02
VI Experimental Results
VI-A Evaluation Metrics
@@
  • Compilability: The proportion of games that can be parsed and compiled by the Ludii game engine. If a game cannot be compiled, it is not evaluated in the functionality metric. The score is normalized from 0 to 100.
  • Functionality: The proportion of playable games. Although a game may compile without errors, it is still considered non-functional if conditions make it unplayable—such as when piece movements rely on undefined positions. The score is normalized from 0 to 100.
@@
  • ROUGE [70]: This is an evaluation metric commonly used in program synthesis to measure the degree of linguistic match between the generated game description and the ground truth game description [71]. It ranges from 0 to 100, with higher values indicating a greater degree of match. The calculation of this metric does not consider syntactic correctness but instead focuses on the similarity between texts. We use the F1 score of ROUGE-L. This value is calculated for each piece of test data, and the average of all data is reported.
  • Normalized Concept Distance (NCD): NCD measures how closely a predicted game matches its ground truth in Ludii. Following [18, 16], games are represented as concept-value vectors derived from semantic features and from behavioral data obtained through automated playouts using a random policy. For example, these values include attributes such as the proportion of turns with at least one legal move and the proportion of the board used at least once. The cosine distance between these vectors gives NCD. Similar to [46], we run 50 playouts for the ground truth and 10 for the predicted games due to computational costs. Since non-functional games cannot compute their concept distance, their distance is set to 1.0. NCD is averaged over all test data, serving as a quality measure for game description generation.
We conducted experiments with 3 different seeds, and each value in the results shows the mean and standard error across the seeds.
VI-B Comparison with Baseline Methods
Table II shows the comparison results with baseline methods. GGDG outperforms GDG with Compilability +37.0, Functionality +30.4, and NCD -0.29, demonstrating that our rule decoding and game description decoding effectively improve the grammatical accuracy of generated game descriptions. In ROUGE, GGDG scores -3.0 lower than GDG. Since ROUGE evaluates the level of reproduction in linguistic expressions, it may undervalue different description methods generated by GGDG that provide the same playable experience. For example, the ending rule for Tic-tac-toe can be defined as either “win by aligning three pieces” or “lose when the opponent aligns three pieces” – while these expressions have the same meaning as rules, they would result in a lower ROUGE score.
While GGDG shows a +4.3 advantage in Compilability compared to SFT+GDG, it performs worse in Functionality (-1.3), ROUGE (-3.5), and NCD (-0.02). This suggests that although GGDG can generate grammatically accurate game descriptions, it still faces challenges compared to SFT in satisfying functionality requirements and reproducing the game quality of the ground truth. SFT+GGDG achieves the best scores across Compilability (72.0), Functionality (70.3), and NCD (0.33) metrics, demonstrating that SFT has a complementary effect on GGDG.
Regarding the baseline results, random generation methods can hardly produce grammatically valid game descriptions. While GDG shows high ROUGE scores, it performs poorly on other metrics. This suggests that although LLMs alone can generate game descriptions that are superficially similar in linguistic expression, without accessing Ludii’s grammar through our decoding methods, they struggle to generate grammatically and functionally accurate game descriptions. Additionally, SFT+GDG shows improved performance across all metrics compared to GDG, indicating that providing the model with Ludii knowledge through SFT further enhances grammatical accuracy.
TABLE III: Ablation study of our proposed method. “Oracle Grammar” indicates cases where grammar $G[y]$ extracted from game description $y$ is used instead of grammar $\hat{G}[y]$ predicted by the LLM. “Grammar” indicates the inclusion of grammar in the query during game description generation. The best results are in bold.
[…]
GDG 27.0$\pm$1.2 26.3$\pm$0.7 63.5$\pm$0.6 0.75$\pm$0.01
GGDG w/o RD, GDD $\checkmark$ 48.3$\pm$0.3 43.3$\pm$1.2 61.6$\pm$0.1 0.59$\pm$0.01
GGDG w/o GDD $\checkmark$ $\checkmark$ 52.0$\pm$3.6 47.3$\pm$2.3 61.3$\pm$0.2 0.55$\pm$0.02
GGDG $\checkmark$ $\checkmark$ $\checkmark$ 64.0$\pm$1.5 56.7$\pm$2.3 60.5$\pm$0.6 0.46$\pm$0.03
Oracle Grammar
GGDG w/o RD, GDD $\checkmark$ 54.0$\pm$1.7 49.3$\pm$0.7 61.7$\pm$0.2 0.53$\pm$0.01
GGDG w/o RD $\checkmark$ $\checkmark$ 62.0$\pm$1.5 57.3$\pm$2.2 64.5$\pm$0.5 0.46$\pm$0.02
VI-C Ablation Study
GGDG Components. We conduct an ablation study on GGDG and summarize the results in Tab. III. We confirm that adding more techniques improves the scores for Compilability, Functionality, and NCD. Among the improvements from GDG to GGDG, grammar accounts for the largest share, comprising 57.6% of Compilability, 55.9% of Functionality, and 55.2% of NCD improvements. This is followed by game description decoding, which accounts for 32.4% of Compilability, 30.9% of Functionality, and 31.0% of NCD improvements, demonstrating their substantial impact. Furthermore, the improvement in Functionality and ROUGE scores when using oracle grammar suggests that there is still room for improvement in grammar generation through rule decoding.